Jake has a before-tax wage of $10 and pays an income tax of 20 percent, so t=.20. Show his budget constraint for work and leisure for one day. Suppose Jake’s before-tax wage decreases to $8, but tax rate remains constant; however, in turns out that Jake still consumes the same amount of leisure. Show this on a graph. Do these results seem reasonable? What do the results tell you about the relative size of the income and substitution effects for leisure for Jake?
Answer #1

Here, if we use an analog to Slutsky Equation, we get:

Total Effect (TE) = Hicksian Substitution Effect (SE) + Income Effect (IE)

\frac{\mathrm{d}L^{M}}{\mathrm{d} w} = \frac{\mathrm{d} L^{H}}{\mathrm{d} w} + ( - L\frac{\mathrm{d} L^{M}}{\mathrm{d} M})

Here, Superscripts M and H stand for Marshallian and Hicksian respectively.

Here, M = Total Income, w = Wage rate, L = Optimal amount of leisure.

per hour Given let L = - Wage w .$.10 Tax rate t 207.b. 2 T = Total time Endowment for a day; Amount of hours of Leisure for